The sum of the ages of Arturo and Brittany is 88 years. 8 years ago, Arturo's age was 3 times Brittany's age. How old is Arturo now?
Answer by lwsshak3(11628)

The sum of the ages of Arturo and Brittany is 88 years. 8 years ago, Arturo's age was 3 times Brittany's age. How old is Arturo now?
let x=Arturo's age
88-x=Brittany'age
8 years ago:
x-8=Arturo's age
(88-x)-8=80-x=Brittany'age
x-8=3(80-x)
x-8=240-3x
4x=248
x=62
88-x=26
Arturo's age now=62